Please solve using LaPlace Partial Fraction expansion for the following Equations (please write very neat, thank you!)

d2y/dt+4dy/dt+3y=30 y(0)=20 y'(0)=12

and

d2y/dt+3dy/dt+2y=24e^-4t y(0)=10 y'(0)=5
Answer #1

* 44 dy Given that: d² dt? dt + 3y = 30 ; y (0)=20 yl(o)=12 Taking Laplace on both sides we have! de 2 +4 24 12 + 3y)=2(30) =بالا ہے 7h Y(0) = 10 +ztp Taking 3dy +2y = 24 e- dt ylcos=5 Laplace on both sides we have a dy + 3 dt2 dt +26 or szy (S) - sy